# Tony Lindeberg

### v1

@@ Abstract
This paper presents an analysis of properties of two hybrid discretization methods for Gaussian derivatives, based on convolutions with either the normalized sampled Gaussian kernel or the integrated Gaussian kernel followed by central differences. The motivation for studying these discretization methods is that in situations when multiple spatial derivatives of different order are needed at the same scale level, they can be computed significantly more efficiently compared to more direct derivative approximations based on explicit convolutions with either sampled Gaussian kernels or integrated Gaussian kernels.
@@
The presented results are intended as a guide, when designing as well as interpreting the experimental results of scale-space algorithms, that operate at very fine scale levels.
@@ 1 Introduction
When to implement the Gaussian derivative operators in scale-space theory, the composed effect of the underlying Gaussian smoothing operation and the following derivative computations need to be discretized in some way. Since a majority of the formulations of scale-space theory are expressed for continuous signals or continuous images, it is essential to also ensure that the desirable properties of the theoretically well-founded scale-space representations are to a sufficiently good degree of approximation transferred to the discrete implementation. Simultaneously, the amount of necessary computations needed for the implementation may often constitute a limiting factor when to choose an appropriate discretization method for expressing actual actual algorithms are to operate on the discrete data to be analyzed.
@@
The subject of this article, is to perform a more detailed analysis of a class of hybrid discretization methods, based on convolution with either the sampled Gaussian kernel, the normalized sampled Gaussian kernel or the integrated Gaussian kernel, followed by computations of discrete derivative approximations by central difference operators, and specifically characterize the degree of approximation of continuous expressions in scale-space theory that these discretization give rise to, see Figure 1 for examples of graphs of equivalent convolution kernels corresponding to these discretization methods. This class of discretization methods was outlined among extensions to future work in Section 6.1 in (Lindeberg 2024), and was also complemented with a description about theoretical scale-space properties of these discretizations in Footnote 13 in (Lindeberg 2024). There were, however, not further in-depth characterizations of the approximation properties of these discretizations with regard to what results they lead to in relation to corresponding results from the continuous scale-space theory.
@@
A further rationale for studying these hybrid discretizations is that in certain applications, such as the use of Gaussian derivative operators in deep learning architectures (Jacobsen et al. 2016, Lindeberg 2021; 2022, Pintea et al. 2021, Sangalli et al. 2022, Penaud-Polge et al. 2022, Gavilima-Pilataxi and Ibarra-Fiallo 2023), the modified Bessel functions of integer order, as used as the underlying mathematical primitives in the discrete analogue of the Gaussian kernel, may, however, not be fully available in the framework used for implementing the image processing operations. For this reason, the hybrid discretizations may, for efficiency reasons, constitute an interesting alternative to using discretizations in terms of either sampled Gaussian derivative kernels or integrated Gaussian derivative kernels, when to implement certain tasks, such as learning of the scale levels by backpropagation, which usually require full availability of the underlying mathematical primitives the scale-parameterized filter family with regard to the deep learning framework, to be able to propagate the gradients between the layers in the deep learning architecture.
@@ 2 Discretization methods for Gaussian derivative operators
    where $I_{n}(s)$ denotes the modified Bessel functions of integer order.
@@
  • the genuinely discrete approach corresponding to convolution with the discrete analogue of the Gaussian kernel $T_{\scriptsize\mbox{disc}}(n;\;s)$ according to (7) followed by central difference operators, thus corresponding to the equivalent discrete approximation kernel (Lindeberg 1993b Equation (58))
@@
To measure how well these different discretization of the Gaussian derivative operators reflect properties of the underlying continuous Gaussian derivatives, let us first consider quantifications in terms of the following spatial spread measure, defined from the absolute value of each equivalent discrete derivative approximation kernel (Lindeberg 2024 Equation (80)):
@@
    As previously stated, the hybrid discretization methods may, however, anyway be warranted in situations where the underlying modified Bessel functions $I_{n}(s)$ are not fully available the computational environment where the discrete filtering operations are to be implemented, such as when performing learning of the scale levels in deep networks based on Gaussian derivative operators coupled in cascade.
@@
For this purpose, we follow a similar methodology as used in (Lindeberg 2024 Section 4). Thus, with now the theory from Section 2 now applied to 2-D image data, by separable application of the 1-D theory along each image dimension, we consider scale-normalized derivative operators defined according to (Lindeberg 1998a; 1998b)
@@ 4.1 Scale-invariant feature detectors with automatic scale selection
    • (i) maxima of the gradient magnitude in the spatial gradient direction $e_{v}$ reformulated such that the second-order directional derivative $L_{vv}$ in the gradient direction is zero and the third-order directional derivative in the gradient direction $L_{vvv}$ is negative (Lindeberg 1998b Equation (8)) and
@@
    $\hat{s}=\operatorname{argmax}_{s}L_{pp,\scriptsize\mbox{norm}}(0,0;\;s)=s_{0},$ (33)
A common property of all these scale-invariant feature detectors is thus the selected scale levels $\hat{s}$ obtained from local extrema over scale will reflect a characteristic scale $s_{0}$ in the input data. By evaluating discretization methods of Gaussian derivatives with respect to such scale selection properties, we therefore have a way of formulating a well-defined proxy task for evaluating how well the different types of discretization methods lead to appropriate consistency properties over scales for the numerical implementation of Gaussian derivative operators.
@@ 4.2 Experimental methodology
To quantify the performance of the different discretization methods, we will
@@
Figure 5: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized Laplacian response according to (21) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,3]$.)
@@
Figure 7: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized determinant of the Hessian response according to (22) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,3]$.)
@@
Figure 9: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized gradient magnitude response according to (26) to a set of diffuse step edges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,3]$.)
@@
Figure 11: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized principal curvature response according to (30) to a set of diffuse ridges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[0.1,3]$.)
@@ 4.3 Experimental results
For generating the input data, we used 50 logarithmically scale values $\sigma_{0}\in[0.1,3]$. When performing the scale selection step, we searched over a range of 80 logarithmically scale levels $\sigma_{0}\in[0.1,5]$, and additionally performed parabolic interpolation over the logarithmic scale values to localized the scale estimates to higher accuracy. This very dense sampling of the scale levels is far beyond what is usually needed in actual computer vision algorithms, but was chosen here, in order to essentially eliminate the effects of sampling issues in the scale direction.
Figure 5 shows a graph of the scale estimates obtained for the second-order Laplacian interest point detector in this way, with the corresponding relative scale errors in Figure 5. Figures 7 and 7 show corresponding results for the non-linear determinant of the second-order Hessian interest point detector.

### v3

@@ Abstract
This paper presents an analysis of properties of two hybrid discretization methods for Gaussian derivatives, based on convolutions with either the normalized sampled Gaussian kernel or the integrated Gaussian kernel followed by central differences. The motivation for studying these discretization methods is that in situations when multiple spatial derivatives of different order are needed at the same scale level, they can be computed significantly more efficiently, compared to more direct derivative approximations based on explicit convolutions with either sampled Gaussian kernels or integrated Gaussian kernels.
@@
The presented results are intended as a guide, when designing as well as interpreting the experimental results of scale-space algorithms that operate at very fine scale levels.
@@ 1 Introduction
When to implement the Gaussian derivative operators in scale-space theory (Koenderink and van Doorn 1987; 1992; Lindeberg 1993a; 1994; 2011; Florack 1997; Weickert et al. 1999; ter Haar Romeny 2003), the composed effect of the underlying Gaussian smoothing operation and the following derivative computations need to be discretized in some way. Since a majority of the formulations of scale-space theory are expressed for continuous signals or continuous images, it is essential to also ensure that the desirable properties of the theoretically well-founded scale-space representations are to a sufficiently good degree of approximation transferred to the discrete implementation. Simultaneously, the amount of necessary computations needed for the implementation may often constitute a limiting factor, when to choose an appropriate discretization method for expressing actual actual algorithms are to operate on the discrete data to be analyzed.
@@
The subject of this article, is to perform a more detailed analysis of a class of hybrid discretization methods, based on convolution with either the sampled Gaussian kernel, the normalized sampled Gaussian kernel or the integrated Gaussian kernel, followed by computations of discrete derivative approximations by central difference operators, and specifically characterize the degree of approximation of continuous expressions in scale-space theory that these discretization give rise to, see Figure 1 for examples of graphs of equivalent convolution kernels corresponding to these discretization methods. This class of discretization methods was outlined among extensions to future work in Section 6.1 in (Lindeberg 2024), and was also complemented with a description about theoretical scale-space properties of these discretizations in Footnote 13 in (Lindeberg 2024). There were, however, no further in-depth characterizations of the approximation properties of these discretizations, with regard to what results they lead to in relation to corresponding results from the continuous scale-space theory.
@@
A further rationale for studying these hybrid discretizations is that in certain applications, such as the use of Gaussian derivative operators in deep learning architectures (Jacobsen et al. 2016, Lindeberg 2021b; 2022, Pintea et al. 2021, Sangalli et al. 2022, Penaud-Polge et al. 2022, Gavilima-Pilataxi and Ibarra-Fiallo 2023), the modified Bessel functions of integer order, as used as the underlying mathematical primitives in the discrete analogue of the Gaussian kernel, may, however, not be fully available in the framework used for implementing the image processing operations. For this reason, the hybrid discretizations may, for efficiency reasons, constitute an interesting alternative to using discretizations in terms of either sampled Gaussian derivative kernels or integrated Gaussian derivative kernels, when to implement certain tasks, such as learning of the scale levels by backpropagation, which usually require full availability of the underlying mathematical primitives the scale-parameterized filter family with regard to the deep learning framework, to be able to propagate the gradients between the layers in the deep learning architecture.
@@ 2 Discretization methods for Gaussian derivative operators
    where $I_{n}(s)$ denotes the modified Bessel functions of integer order (Abramowitz and Stegun 1964).
@@
  • the genuinely discrete approach corresponding to convolution with the discrete analogue of the Gaussian kernel $T_{\scriptsize\mbox{disc}}(n;\;s)$ according to (7) followed by central difference operators $\delta_{x^{\alpha}}$, thus corresponding to the equivalent discrete approximation kernel (Lindeberg 1993b Equation (58))
@@
To measure how well these different discretization of the Gaussian derivative operators reflect properties of the underlying continuous Gaussian derivatives, let us consider quantifications in terms of the following spatial spread measure, defined from the absolute value of each equivalent discrete derivative approximation kernel (Lindeberg 2024 Equation (80)):
@@
    As previously stated, the hybrid discretization methods may, however, anyway be warranted in situations where the underlying modified Bessel functions $I_{n}(s)$ are not fully available in the computational environment where the discrete filtering operations are to be implemented, such as when performing learning of the scale levels in deep networks based on Gaussian derivative operators coupled in cascade.
@@
For this purpose, we follow a similar methodology as used in (Lindeberg 2024 Section 4). Thus, with the theory in Section 2 now applied to 2-D image data, by separable application of the 1-D theory along each image dimension, we consider scale-normalized derivative operators defined according to (Lindeberg 1998a; 1998b)
@@ 4.1 Scale-invariant feature detectors with automatic scale selection
    • (i) maxima of the gradient magnitude in the spatial gradient direction $e_{v}$, reformulated such that the second-order directional derivative $L_{vv}$ in the gradient direction is zero and the third-order directional derivative in the gradient direction $L_{vvv}$ is negative (Lindeberg 1998b Equation (8)) and
@@
    $\hat{s}=\operatorname{argmax}_{s}L_{pp,\scriptsize\mbox{norm}}(0,0;\;s)=s_{0}.$ (33)
A common property of all these scale-invariant feature detectors is thus the selected scale levels $\hat{s}$ obtained from local extrema over scale will reflect a characteristic scale $s_{0}$ in the input data (Lindeberg 2021a). By evaluating discretization methods of Gaussian derivatives with respect to such scale selection properties, we therefore have a way of formulating a well-defined proxy task for evaluating how well the different types of discretization methods lead to appropriate consistency properties over scales for the numerical implementation of Gaussian derivative operators.
@@ 4.2 Experimental methodology
To quantify the performance of the different discretization methods with regard to scale selection tasks, we will
@@
Figure 5: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized Laplacian response according to (21) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,3]$.)
@@
Figure 7: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized determinant of the Hessian response according to (22) to a set of Gaussian blobs of different size $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,3]$.)
@@
Figure 9: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized gradient magnitude response according to (26) to a set of diffuse step edges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,3]$.)
@@
Figure 11: Graphs of the relative scale estimation error $E_{\scriptsize\mbox{scaleest,rel}}(\sigma)$, according to (34), when applying scale selection from local extrema over scale of the scale-normalized principal curvature response according to (30) to a set of diffuse ridges of different width $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}$, for different discrete approximations of the Gaussian derivative kernels, for either discrete analogues of Gaussian derivative kernels $T_{\scriptsize\mbox{disc},x^{\alpha}}(n;\;s)$ according to (10), sampled Gaussian derivative kernels $T_{\scriptsize\mbox{sampl},x^{\alpha}}(n;\,s)$ according to (8), integrated Gaussian derivative kernels $T_{\scriptsize\mbox{int},x^{\alpha}}(n;\,s)$ according to (9), the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-sampl},x^{\alpha}}(n;\;s)$ according to (16) or the hybrid discretization method corresponding the equivalent convolution kernels $T_{\scriptsize\mbox{hybr-int},x^{\alpha}}(n;\;s)$ according to (17). (Horizontal axis: Reference scale $\sigma_{\scriptsize\mbox{ref}}=\sigma_{0}\in[1/3,3]$.)
@@ 4.3 Experimental results
For generating the input data, we used 50 logarithmically scale values $\sigma_{0}\in[0.1,3]$. When performing the scale selection step, we searched over a range of 80 logarithmically scale levels $\sigma_{0}\in[0.1,5]$, and accumulated the scale-space signature over scale at the image center for each differential feature detector, and additionally performed parabolic interpolation over the logarithmic scale values to localize the scale estimates to higher accuracy. This very dense sampling of the scale levels is far beyond what is usually needed in actual computer vision algorithms, but was chosen here, in order to essentially eliminate the effects of discrete sampling issues in the scale direction.
Figure 5 shows a graph of the scale estimates obtained for the second-order Laplacian interest point detector in this way, with the corresponding relative scale errors in Figure 5. Figures 7 and 7 show corresponding results for the non-linear determinant of the Hessian interest point detector.